The following budget information relates to General Motors Company.

Variable manufacturing overhead costs per machine hour are as follows

Expected annual operating capacity               120,000 Direct Labor Hours

Variable overhead costs

     Indirect labor                                        $310,000

     Indirect materials                                     100,000

     Factory supplies                                        34,000

         Total variable                                  444,000

Fixed overhead costs

     Depreciation                                           170,000

     Supervision                                            120,000

     Property taxes                                          96,000

         Total fixed                                     386,000

Total costs       $830,000

The relevant range for monthly activity is expected to be between 9,000 and 11,000 direct labor hours.

Instructions

General Motors would like a flexible manufacturing overhead budget

prepared based on machine hours. Given the range of activity above,

Prepare the budget using increments of 1,000 machine hours.
Answer #1

Working notes:

Computation of variable overhead cost per machine hour:

Variable overhead cost per machine hour = variable overhead cost / total expected annual machine hours

Indirect labor = $310,000 / 120,000 = $2.583

Indirect material = $100,000 / 120,000 = $0.833

Factory supplies = $34,000 / 120,000 = $0.283

General Motors

Flexible Manufacturing Overhead Budget

Activity level
Machine hours 9,000 10,000 11,000
Variable cost:
Indirect labor ($2.583) $ 23,247 $ 25,830 $ 28,413
Indirect material ($0.833) 7,497 8,330 9,163
Factory supplies ($0.283) 2,547 2,830 3,113
Total variable cost 33,291 36,990 40,689
Fixed cost:
Depreciation ($170,000 / 12) 14,167 14,167 14,167
Supervision ($120,000 / 12) 10,000 10,000 10,000
Property taxes ($96,000 / 12) 8,000 8,000 8,000
Total fixed cost 32,167 32,167 32,167
Total cost $ 65,458 $ 69,157 $ 72,856